From GlueXWiki
Jump to: navigation, search

July 18, 2013 FDC meeting


  1. Production
    • Final package assembly [1], [2], [3] (Bill, Dave)
  2. Engineering (Bill)
    • Survey
    • Connector strain relief [4], [5],[6]
  3. Electronics (Chris, Nick)
  4. Full electronics test FDC E-log (Lubomir, Vlad)
  5. Other


Participants: Bill, Chris, Nick, Simon, Fernando, Vlad, Eugene, Glenn, and Lubomir.


- On Monday we finished the tests with the third package and it was moved to Blue Crab. On the same day the spare package was moved to JLab.

- Final package installation (see the pictures linked above): All the four packages were connected with the spacers. Like between with the first spacer we had to put additional rod in the second spacer to clear the gap between the spacer and the gusset ring. The last (shorter) spacer didn't need additional threaded rod, at least visually.


- Survey: Jim Dahlberg was at Blue Crab yesterday and will start surveying the packages possibly tomorrow using 12ft articulated arm. The idea is to check (and correct) the flatness of each package and how parallel they are.

- Based on our experience with the test with full electronics we need good strain relief for the connectors. Therefore, Bill decided to make additional skim above the spacers to be used to attach the cables (see the model details linked above). It will use 15mil g10 (or lighter material) and be supported by ~22 or less plastic (or g10) holders attached to the rods. There will be cuts on the skin to insert cable ties. Bill will start prototyping it very soon.


- There were long discussions about the grounding of the packages. We will have one ground cable clamped to all the packages at the holding brackets and possibly a separate removable ground cable connecting the packages to the clean ground. There will be some additional ground connections to other parts inside the package, but using thinner cables. The cooling manifolds are connected to the same ground. Fernando will stop at Blue Crab to make sure that grounding is done correctly and what modifications are needed eventually.

- Nick and Chris will fix the setting on the VXS power supply in 126. Fernando: this will further reduce the noise on the fADCs as it happened with the fADC250s (2.4kHz effect).

- Nick on the spare pre-amps: 14 pre-amps cathode type and 8 pre-amps anode type.

Full electronics tests

- The problematic card on D5 (channels 97A-120) with ~10 bad channels was fixed by inserting a thin (15mil) g10 plate in between the cathode and wire frame, thus applying pressure on the conductive tape. The plate was sticking out of the frame pushing the end of the pre-amp down. In this position there were 2 bad channels (Elog entry 116), but after Casey cut the plate and the pre-amp return to its normal position all the channels were working. However two channels of the next pre-amp stopped working.

- All the rest of the channels were fixed by touching the card or bending the cooling tube. In one case it was a wire pre-amp showing the same problems. Therefore, it is not only the conductive tape that causes these problems.

- We took data over the last weekend also at two other HV. For these two runs the F1TDC crate showed some problems that have to be understood.

- Now we have the spare package at 126 being flushed with gas.


- No meetings next two week. Next meeting on 8/8.